# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v1

@@ 2.4 Fino1: LLMs for Financial Reasoning
For reasoning path generation, we utilize GPT-4o, as it demonstrates consistently strong performance across multiple datasets (as reported in Table 3), achieving results comparable to the highest-performing models on FinQA. Followed by HuatuoGPT-o1 Chen et al. (2024), we adopt a similar strategy to refine the CoT reasoning paths generated by GPT-4o. The refinement process involves the use of a verifier to assess the correctness of the generated CoT. If the verifier rejects the reasoning path, the model iteratively refines it by applying a strategy sampled from one of the following: (1) Backtracking – Revisiting previous steps to correct logical inconsistencies. (2) Exploring New Paths – Generating alternative reasoning routes when the current path fails. (3) Verification – Cross-checking intermediate steps to ensure logical soundness. (4) Correction – Making targeted adjustments to rectify errors. The CoT refinement process begins with an initial prompt to generate a reasoning path and an answer. The verifier then evaluates the response. If the answer is incorrect, one of the above strategies is applied iteratively to refine the CoT process until a correct and logically sound answer is generated.
2.4.1 Mode Training
After curating the reasoning data, we fine-tune Llama3.1-8B-Instruct, selected as the optimal backbone model based on the performance of 8B models. First, the curated dataset is used for supervised fine-tuning (SFT), where the question serves as input, and the reasoning path along with the final answer serves as output, enabling the model to learn structured financial reasoning. Next, we apply reinforcement learning with a reward model trained to assess the quality of generated reasoning paths. Inspired by prior work, we employ a verifier-based sparse reward mechanism and optimize the model using Proximal Policy Optimization (PPO). The verifier evaluates the correctness and coherence of the reasoning path; if an error is detected, the model iteratively refines the reasoning steps through backtracking, exploring alternative paths, verification, and correction. This two-stage approach enhances logical consistency, financial reasoning depth, and equation-solving accuracy, ultimately improving model performance across financial benchmarks.
During the SFT phase, the model was trained with a learning rate of 5e-6 for three epochs, using a maximum sequence length of 8192 tokens. LoRA (Low-Rank Adaptation) was applied to update parameters efficiently, reducing computational costs while maintaining performance. In the subsequent RL stage, the learning rate was lowered to 5e-7, and the model was trained for an additional three epochs. The reward model, previously trained in Chen et al. (2024), was employed to evaluate the quality of the model’s reasoning paths. This reward signal guided optimization through Proximal Policy Optimization (PPO), refining the model’s ability to generate accurate and coherent financial reasoning.
@@ 3 Results
3.1 Overall Results Analysis
The resutls of all these models on three dataset are listed in Table 3. Among all models, DeepSeek-R1 achieves the highest overall performance (68.93), driven by its strong XBRL-Math results, with its distilled variants, DeepSeek-R1-Distill-Llama-70B and DeepSeek-R1-Distill-Qwen-32B, closely following. Among large-scale models, GPT-4o performs well across datasets but is outperformed by DeepSeek-R1 due to its lower XBRL-Math score. General-purpose 70B models, such as Qwen2.5-72B-Instruct and Llama3.3-70B-Instruct, achieve strong results, with Llama3.3-70B showing significant improvement over its predecessors. Surprisingly, these models also surpass the reasoning-focused GPT-o1 and approach GPT-o3-mini, suggesting that general-purpose reasoning strategies do not always translate effectively to financial tasks. In the 8B category, Fino1-8B (61.03) outperforms other models, demonstrating robust task optimization. Notably, reasoning-enhanced models exhibit mixed results: while they improve XBRL-Math scores, they tend to perform worse on FinQA and DM-Simplong. This suggests that different reasoning strategies yield varying benefits across financial reasoning tasks, emphasizing the importance of task-specific optimization.
Models FinQA DM-Simplong XBRL-Math Average
GPT-4o 72.49 60.00 72.22 68.24
GPT-o1 49.07 56.00 74.44 59.84
GPT-o3-mini 60.87 59.00 76.67 65.51
DeepSeek-V3 73.20 53.00 76.67 67.62
DeepSeek-R1 65.13 53.00 86.67 68.93
Qwen2.5-72B-Instruct 73.38 59.00 67.78 66.72
Qwen2.5-72B-Instruct-Math 69.74 42.00 83.33 65.69
DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 68.80
Llama3-70B-Instruct 58.92 41.00 56.67 52.20
Llama3.1-70B-Instruct 63.18 48.00 63.33 58.17
Llama3.3-70B-Instruct 68.15 54.00 70.00 64.05
DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 68.97
DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 63.90
DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 53.36
Llama3-8B-Instruct 41.97 29.00 48.89 39.95
Llama3.1-8B-Instruct 54.13 34.00 62.22 50.12
Fino1-8B 60.87 40.00 82.22 61.03
@@ 3.1 Overall Results Analysis
Table 3: Performance of different LLMs on three tested financial datasets.
From our results, we identify several key insights regarding the performance of large language models (LLMs) in financial reasoning tasks:
(1) General reasoning enhancements do not necessarily translate to improvements in financial tasks. While models fine-tuned for general reasoning, such as GPT-o1 and DeepSeek-R1, demonstrate strong performance in mathematical and logical reasoning tasks, they do not consistently outperform other models in financial benchmarks. For example, GPT-o1 scores 49.07 in FinQA and 56.0 in DM-Simplong, trailing behind models like GPT-4o that are not explicitly optimized for general reasoning. Also, DeepSeek-R1 also under performan V3 in these tasks. This suggests that financial reasoning requires more than just enhanced logical and mathematical capabilities; it demands a deep understanding of domain-specific terminology, structured financial statements, and numerical patterns. One possible explanation for this phenomenon is that reasoning-focused fine-tuning, which emphasizes general problem-solving strategies, may inadvertently weaken a model’s retention of financial knowledge. This is particularly relevant for tasks requiring precise comprehension of financial terms, balance sheet structures, and industry-specific metrics, where models without targeted financial fine-tuning may struggle. Additionally, reasoning-enhanced models might prioritize general logical inference over domain-specific factual recall, which could lead to a decrease in performance when interpreting specialized financial contexts. These findings indicate that while general reasoning improvements enhance broad problem-solving skills, they do not necessarily translate into domain-specific advantages, highlighting the need for specialized financial training in LLMs.
(2) Different reasoning enhancement strategies yield varying improvements depending on the financial reasoning task format. By comparing the performance of GPT-4o vs. GPT-o1, Qwen2.5-72B-Instruct vs. Qwen2.5-72B-Instruct-Math, and Llama models vs. the DeepSeek-R1 distilled Llama models, we observe that reasoning-enhanced models do not always improve financial task performance. GPT-o1 underperforms GPT-4o in FinQA (49.07 vs. 72.49) and DM-Simplong (56.00 vs. 60.00), suggesting that its reasoning strategies, such as chain of thought and reflection-tuning, are optimized for general problem-solving rather than financial-specific logic. Similarly, Qwen2.5-72B-Instruct-Math, despite achieving a substantial XBRL-Math boost (67.78 → 83.33), experiences sharp declines in FinQA (69.74 → 42.00) and DM-Simplong (59.00 → 42.00), indicating that mathematical reasoning enhancements help structured financial data but do not generalize well to broader financial tasks. In the 8B category, DeepSeek-R1-Distill-Llama-8B (53.36) outperforms Llama3.1-8B-Instruct (50.12) and Llama3-8B-Instruct (39.95), excelling in XBRL-Math (81.11) but showing limited improvements in other financial tasks.
These results suggest that different reasoning strategies target distinct aspects of financial reasoning. DeepSeek-R1 and Qwen2.5-72B-Instruct-Math demonstrate strong performance in symbolic and mathematical reasoning, excelling in structured equation-based tasks. However, they do not generalize well to financial question-answering tasks like FinQA and DM-Simplong. Similarly, reasoning augmentations in GPT-o1 fail to provide consistent improvements, likely due to the absence of relevant financial reasoning data, highlighting the challenge of adapting broad reasoning techniques to financial logic.
(3) Scaling model size does not always lead to performance gains in financial tasks. While larger models typically achieve better results in standard NLP tasks, their performance in financial tasks does not always scale proportionally with size. For instance, DeepSeek-70B (66.73 in FinQA, 53 in DM-Simplong, and 86.67 in XBRL-Math) performs similarly to its smaller distilled variant DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite the latter having less than half the parameters. A broader comparison, including Qwen2.5-72B-Instruct, DeepSeek-R1-Distill-Llama-70B, GPT-4o, DeepSeek-V3, and DeepSeek-R1, reveals that performance stabilizes once model size reaches approximately 70B parameters. One possible explanation for this trend is the nature of XBRL-Math, which consists of structured financial tables and numerical equations that prioritize precise mathematical computations over complex language-based reasoning. In such structured, numerically driven tasks, smaller models can achieve competitive performance by leveraging arithmetic operations and pattern recognition rather than advanced language modeling. This suggests that the advantage of larger models diminishes in financial decision-making tasks where structured data comprehension and numerical reasoning play a more crucial role than language-based inference.
(4) Improving Pre-training Data and Strategies in General Domains Also Benefits Financial Tasks. Compared with Llama 3-8B-Instruct, Llama 3.1-8B-Instruct demonstrates a significant performance boost, increasing its average score from 39.95 to 50.12, with substantial improvements in XBRL-Math (62.22 vs. 48.89) and FinQA (54.13 vs. 41.97). Similarly, among the 70B-Instruct models, Llama 3.1-70B-Instruct outperforms Llama 3-70B-Instruct (58.17 vs. 52.20), with the largest improvement in DM-Simplong (48.00 vs. 41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05 vs. 58.17), excelling in FinQA (68.15 vs. 63.18) and XBRL-Math (70.00 vs. 63.33), highlighting the benefits of dataset expansion and refined training strategies.
More specifically, Llama 3.1 incorporates a more extensive dataset, covering over 15 trillion tokens across multiple languages and domains, leading to notable performance gains. Llama 3.3 further benefits from advancements in post-training techniques, such as online preference optimization, which enhances core performance while significantly reducing training and inference costs. These refined optimization strategies in pre-training not only improve general capabilities but also enhance performance in domain-specific tasks.
(5) Existing reasoning models struggle with long financial contexts and cross-table correlations. Many financial reasoning tasks require models to extract and synthesize information from lengthy reports, structured tables, and numerical equations. However, performance discrepancies across datasets suggest that current LLMs struggle with this challenge. For example, models that perform well in FinQA, which primarily involves textual financial reasoning, do not necessarily excel in DM-Simplong, which integrates numerical and symbolic information. GPT-4o achieves 72.49 in FinQA but drops to 60.0 in DM-Simplong, indicating difficulties in integrating mathematical symbols with textual reasoning. Furthermore, models that score highly in XBRL-Math, such as DeepSeek-R1-Distill-Qwen-32B (84.44) and DeepSeek-R1 (86.67), highlight a growing but incomplete capability to process structured financial tables. Notably, all reasoning-enhanced models fail to show improvements in handling long financial contexts and multi-table correlations. This is particularly concerning given that financial reports are inherently long and include complex interlinked tables. Though smaller models perform well in structured inputs like XBRL-Math, their performance drops significantly—by nearly 30%—when handling long contexts. For example, DeepSeek-R1-Distill-Llama-8B achieves 81.11 in XBRL-Math but drops to 33.0 in DM-Simplong. This suggests that while smaller models can recognize structured patterns, they struggle with long-range dependencies and multi-step reasoning. Addressing these limitations requires improvements in long-context retention, structured data comprehension, and multi-modal document understanding to enhance financial reasoning capabilities.
(6) Reasoning enhancement basd on financial data improve performance of all datasets. Compared with Llama 3.1-8B-Instruct, DeepSeek-R1-Distill-Llama-8B, which incorporates reasoning enhancement data distilled from DeepSeek-R1, shows only a modest improvement (50.12 to 53.36). However, our model, Fino1-8B, trained with reasoning paths distilled from GPT-4o using a backtracking strategy to iteratively refine the reasoning process, achieves a substantial performance boost across all datasets, with an average score of 61.03—a 10.91-point increase over Llama 3.1-8B-Instruct. Notably, Fino1-8B outperforms its counterparts in FinQA (60.87 vs. 54.13), DM-Simplong (40.00 vs. 34.00), and particularly in XBRL-Math (82.22 vs. 62.22), demonstrating the effectiveness of reasoning-driven training.
It is important to note that our training utilized reasoning paths generated exclusively from FinQA, yet the model exhibited improvements across all datasets. This suggests that financial domain-specific reasoning enhancement data can significantly improve models’ comprehensive understanding of financial terminology, long-context comprehension, table-based reasoning, and equation interpretation. Furthermore, these results highlight the effectiveness of a combination of reasoning path-building strategies, including backtracking, exploring new paths, verification, and correction, in financial tasks. By refining reasoning paths through iterative corrections, these strategies enhance logical consistency, computational accuracy, and overall robustness, demonstrating a promising approach for advancing financial reasoning capabilities in language models.
@@
Overall, these insights demonstrate that financial reasoning presents unique challenges distinct from general NLP and mathematical reasoning. The variability in performance across different tasks underscores the necessity for more financially specialized training methodologies and architectural improvements to enhance financial comprehension, structured data processing, and multi-step numerical reasoning.
3.2 Error Analysis

### v2

@@ 2.5 Reasoning Path Generation and Refinement
For reasoning path generation, we use GPT-4o, chosen for its strong performance across datasets (Table 3) and its stability in generating high-quality outputs. Following the CoT refinement framework from HuatuoGPT-o1 Chen et al. (2024), we employ a verifier-based iterative process to refine the reasoning paths generated by GPT-4o. If a reasoning path is rejected by the verifier, refinement is guided by one of the following strategies: (1) Backtracking: Revisiting previous steps to correct logical inconsistencies. (2) Exploring New Paths: Generating alternative reasoning routes when the current path fails. (3) Verification: Cross-checking intermediate steps to ensure logical soundness. (4) Correction: Making targeted adjustments to rectify specific errors. The CoT refinement process begins with an initial prompt to generate a reasoning path and final answer. If the verifier identifies the answer as incorrect, one of the above strategies is applied iteratively until a correct and logically consistent answer is produced.
2.5.1 Model Training
We develop our financial reasoning models based on curated reasoning paths, focusing on smaller-sized models for efficient real-world deployment. We select LLaMA-3.1-8B-Instruct and Qwen2.5-14B-Instruct as backbone models, given their strong performance among small and mid-sized LLMs. We adopt a two-stage training approach. In the first stage, we perform Supervised Fine-Tuning (SFT), using questions as input and reasoning paths with final answers as output. For the 8B model, we use only FinQA reasoning paths to assess the impact of domain-specific fine-tuning. For the 14B model, we use the full set of sampled and filtered data to leverage its greater capacity for diverse and complex reasoning. In the second stage, we apply RL using a verifier-based sparse reward mechanism and optimize with Proximal Policy Optimization (PPO), following Chen et al. (2024). A reward model evaluates the correctness and coherence of the reasoning path. If errors are found, the model iteratively refines its outputs through backtracking, alternative path exploration, verification, and correction, improving logical consistency and equation-solving accuracy. During SFT, models are trained for three epochs with a learning rate of 5e-6 and a maximum sequence length of 8192 tokens. LoRA is applied for efficient parameter updates. In the RL stage, training continues for three epochs with a reduced learning rate of 5e-7, guided by the reward model to refine reasoning quality via PPO.
@@ 3 Results
Table 3 presents the performance of various LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong. Our models, Fino1-8B and Fino1-14B, demonstrate strong and consistent results across tasks, outperforming several larger and closed-source models. Fino1-14B achieves an average score of 60.25, surpassing many larger-scale models such as LLaMA-3.3-70B-Instruct (56.04), Qwen2.5-72B-Instruct (53.71), GPT-o1-preview (54.05), GPT-o3-mini (57.89), and even GPT-4.5 (60.43). It also achieves the best performance among all models on two benchmarks—DM-Simplong (60.00) and XBRL-Math (86.67). These results demonstrate that high-quality, domain-specific reasoning training can enable smaller models to rival—and even surpass—the performance of significantly larger and proprietary LLMs.
Fino1-8B also performs competitively with an average score of 50.77, surpassing other 8B models and even some 14B and 32B models. Both models excel on XBRL-Math, where Fino1-14B achieves the top score of 86.67. Moreover, Fino1-14B and Fino1-8B significantly outperform FinR1-7B, a financial reasoning-enhanced model trained on a much larger 60K task-aligned dataset, demonstrating that larger data volume does not necessarily lead to better performance, and that carefully curated, high-quality reasoning data—even in smaller quantities—can yield superior results.
Models FinQA DM-Simplong XBRL-Math DM-Complong Average
GPT-4o 72.49 60.00 72.22 39.33 61.01
GPT-o1-preview 49.07 56.00 74.44 36.67 54.05
GPT-o3-mini 60.87 59.00 76.67 35.00 57.89
DeepSeek-V3 73.20 53.00 76.67 42.33 61.30
DeepSeek-R1 65.13 53.00 86.67 38.67 60.87
GPT-4.5 68.94 59.00 74.44 39.33 60.43
Meta-Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
Llama-3.1-70B-Instruct 63.18 48.00 63.33 34.33 52.21
Llama-3.3-70B-Instruct 68.15 54.00 70.00 32.00 56.04
Qwen2.5-72B-Instruct 73.38 59.00 67.78 14.67 53.71
Qwen2.5-Math-72B-Instruct 69.74 42.00 83.33 5.00 50.02
DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 30.67 59.27
Qwen2.5-32B-Instruct 73.11 56.00 65.56 30.00 56.17
Qwen/QwQ-32B 61.22 46.00 84.44 20.00 52.92
DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 24.67 57.40
Limo 63.44 45.00 61.11 15.33 46.22
S1-32B 66.81 53.00 84.44 24.00 57.06
Qwen2.5-14B-Instruct 67.44 59.00 57.78 26.67 52.72
DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 21.00 53.18
DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 15.67 43.94
Meta-Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
Llama-3.1-8B-Instruct 54.13 34.00 62.22 14.30 41.16
Qwen2.5-7B-Instruct 55.37 41.00 42.22 17.67 39.07
FinR1-7B 58.74 37.00 30.00 13.67 34.85
Fino1-8B 60.87 40.00 82.22 20.00 50.77
Fino1-14B 70.01 60.00 86.67 24.33 60.25
@@
Table 3: Performance of different LLMs on four financial reasoning datasets.
Beyond overall performance, the comparison across models provides additional insights that can guide the development of LLMs for financial applications:
(1) General reasoning enhancements do not necessarily improve financial task performance. Models like GPT-o1 and DeepSeek-R1, though strong in general reasoning, underperform on financial benchmarks, highlighting the importance of domain-specific knowledge and structured numerical reasoning. Moreover, reasoning enhancement strategies based on general-domain data, as seen in S1-32B and Limo, often fail to improve and sometimes even degrade performance compared to their backbone models (e.g., Qwen2.5-32B). These results underscore that high-quality, domain-aligned reasoning paths are more effective than general reasoning augmentation, and tailored financial training is crucial for optimal model performance.
(2) Reasoning enhancements vary in effectiveness across financial tasks. Models like GPT-o1 perform worse than GPT-4o on FinQA (49.07 vs. 72.49) and DM-Simplong (56.00 vs. 60.00). Qwen2.5-Math improves in XBRL-Math (67.78 $\rightarrow$ 83.33) but drops sharply in FinQA (69.74 $\rightarrow$ 42.00) and DM-Simplong (59.00 $\rightarrow$ 42.00). DeepSeek-R1-Distill-Llama-8B also excels in XBRL-Math (81.11) but shows limited gains elsewhere. These results suggest that strategies like CoT or math fine-tuning may offer benefits for structured tasks but often fail to generalize effectively to broader financial reasoning. Compared with Qwen2.5-72B-Instruct and its math enhanced version, we can see that while TIR and synthetic building appears to help in equation-heavy tasks like XBRL-Math, its impact on other types of financial tasks is less clear and may be influenced by multiple factors such as task structure, data modality, or domain complexity.
(3) Larger model size offers limited benefits in financial reasoning. While larger models typically excel in general NLP tasks, their gains in financial tasks are not proportional. For example, DeepSeek-70B (66.73 FinQA, 53 DM-Simplong, 86.67 XBRL-Math) performs similarly to the smaller DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite having less than half the parameters. Across models like Qwen2.5-72B, DeepSeek-R1-Distill-Llama-70B, and GPT-4o, performance tends to plateau around 70B. This may reflect the structured nature of tasks like XBRL-Math, where arithmetic precision outweighs complex language reasoning, enabling smaller models to compete effectively.
(4) Improved pre-training data and strategies in general domains also benefit financial tasks. Llama 3.1-8B-Instruct shows a notable performance boost over Llama 3-8B-Instruct (avg. 50.12 vs. 39.95), with gains in XBRL-Math (62.22 vs. 48.89) and FinQA (54.13 vs. 41.97). Similarly, Llama 3.1-70B-Instruct outperforms Llama 3-70B (58.17 vs. 52.20), especially in DM-Simplong (48.00 vs. 41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05), driven by improvements in FinQA (68.15) and XBRL-Math (70.00). These gains stem from larger, more diverse training corpora (e.g., 15T+ tokens in Llama 3.1) and enhanced post-training techniques like online preference optimization. Such strategies improve general reasoning while also boosting performance in domain-specific tasks like financial reasoning.
(5) Existing reasoning models struggle with long contexts and cross-table correlations. Financial tasks often require integrating information from lengthy reports, structured tables, and equations, yet current models show clear limitations. For example, GPT-4o performs well on FinQA (72.49) but drops on DM-Simplong (60.00), indicating challenges in combining textual and symbolic reasoning. Models like DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-32B excel in XBRL-Math but fail to generalize. Performance on DM-Complong remains low across models (e.g., LLaMA-3.3-70B: 32.00). Smaller models like DeepSeek-R1-Distill-Llama-8B also perform poorly with 33.00 for DM-Simplong. These results highlight the need for improved long-context modeling and structured data integration in financial reasoning.
(6) Financial reasoning enhancement improves performance across all tasks. Compared to LLaMA 3.1-8B-Instruct (41.16), DeepSeek-R1-Distill-Llama-8B shows a modest gain (43.94). In contrast, our Fino1-8B, trained on GPT-4o-distilled reasoning paths with backtracking strategies, achieves a substantial improvement, reaching an average score of 50.77, and consistently outperforms baselines on FinQA (60.87 vs. 54.13), DM-Simplong (40.00 vs. 34.00), and XBRL-Math (82.22 vs. 62.22). Fino1-14B further boosts performance, achieving an average score of 60.25, outperforming other 14B models (e.g., Qwen2.5-14B: 52.72) and even rivaling state-of-the-art models like DeepSeek-R1 (60.87) and GPT-4o (61.01). Notably, although Fino1-8B and Fino1-14B were trained with limited financial CoTs, both models show improvements across all tasks, highlighting the effectiveness of high-quality reasoning paths and structured training strategies such as backtracking, path exploration, and verification.
@@
Overall, these insights highlight that financial reasoning poses unique challenges distinct from general NLP and mathematical reasoning. The performance variability across tasks underscores the need for specialized training strategies and architectural improvements to better support financial comprehension, structured data processing, and multi-step numerical reasoning.
4 Conclusion
This study presents a comprehensive evaluation of 24 general-purpose and reasoning-augmented LLMs on financial reasoning tasks, encompassing financial text understanding, numerical computation, and structured data interpretation. Our findings show that existing models, despite strong performance on general reasoning benchmarks, often fail to generalize to financial domains—particularly in handling long contexts, multi-table reasoning, and domain-specific terminology. To bridge this gap, we curated a high-quality, diverse financial reasoning dataset and introduced Fino1, a suite of domain-adapted models trained with GPT-4o-generated reasoning paths using chain-of-thought fine-tuning and reinforcement learning. Despite being trained on limited data, Fino1 models outperform several larger and advanced LLMs, including GPT-o1, GPT-o3-mini, GPT-4.5, and DeepSeek models (V3 and R1). These results highlight the effectiveness of domain-specific reasoning strategies and demonstrate the practical value of tailored model adaptation for real-world financial applications.
Appendix A Related Work
A.1 General-Purpose and Reasoning-Enhanced LLMs
A.2 LLMs in Financial Applications
A.3 Reasoning Capabilities in Financial Tasks
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al. (2021), mathematical problem-solving in financial contexts Srivastava et al. (2024), and structured data interpretation Liu et al. (2025a). These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
Appendix B Details Data Source
We provide here detailed descriptions of the datasets used to construct our training and evaluation corpus for financial reasoning enhancement.
  • FinQA Chen et al. (2022): FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
  • DocFinQA Reddy et al. (2024b): An extension of FinQA designed for long-context reasoning. It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words. This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
  • TAT-QA Zhu et al. (2021): A large-scale financial QA dataset created from real financial reports. It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison. TAT-QA reflects real-world challenges in financial document interpretation.
  • DocMath-Eval Zhao et al. (2024b): This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
    • DocMath-CompLong: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
    • DocMath-SimpShort: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
  • BizBench-QA Krumdick et al. (2024): Contains 14,400 financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode. These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
Figure 1: Error case 1.
Figure 2: Error case 2.
Appendix C Error Analysis